OpenStudy (anonymous):

If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4

OpenStudy (dumbcow):

it takes 4.062 seconds for the ball to hit the ground

OpenStudy (anonymous):

that isnt an option

OpenStudy (anonymous):

\[2+\left(\begin{matrix}\sqrt{}17 \\ 2\end{matrix}\right)\]